OpenStudy (aravindg):

if f(X)=sqrt(x-3) g(x)=(x^2-2x+1)/(x^2-8x+12) find (f+g)x anf fg

OpenStudy (aravindg):

today i hav exam ...

OpenStudy (anonymous):

first one just add them up

OpenStudy (anonymous):

\[(f+g)(x)=f(x)+g(x)\]

OpenStudy (aravindg):

show plzz i think adding is complicate

OpenStudy (anonymous):

write them both down a stick a plus sign between them

OpenStudy (aravindg):

no need to add??

OpenStudy (anonymous):

\[\sqrt{x-3} +\frac{x^2-2x+1}{x^2-8x+12}\]

OpenStudy (aravindg):

only that much needed??what about fg?

OpenStudy (anonymous):

you feel like adding that? it will be a big stupid mess and will look even worse if you put it over one denominator.

OpenStudy (anonymous):

\[fg(x)=f(x)g(x)\]

OpenStudy (anonymous):

put them in parentheses next to each other

OpenStudy (anonymous):

i.e. replace the plus sign above by a times sign. finished

OpenStudy (aravindg):

wow thx

OpenStudy (anonymous):

i wish all math was this easy right?

OpenStudy (aravindg):

what u mean by general solution of cos x=1/2

OpenStudy (anonymous):

cosine is a periodic function, that is \[\cos(x+2\pi)=\cos(x)\] so there are an infinite number of solutions to \[\cos(x)=\frac{1}{2}\]

OpenStudy (anonymous):

"general solution" means find some expression for all of them, rather than just one answer